OpenStudy (anonymous):

Sam Monte deposits $21,500 into Legal Bank which pays 6 percent interest that is compounded semiannually. what will Sam have in his account at the end of 6 years?

OpenStudy (dcolley):

30,653.86

OpenStudy (anonymous):

\[A = P(1+r)^{t}\]where P is the principal, r is the rate (over the compounding period, expressed as a decimal), and t is the number of periods to compound over.

OpenStudy (dcolley):

its A=P(1+/n)^nt

OpenStudy (dcolley):

I meant A=P(1+r/n)^nt

OpenStudy (anonymous):

http://openstudy.com/groups/mathematics/updates/4e8fbb4f0b8bfb9f23db7726#/groups/mathematics/updates/4e8fc48f0b8bfb9f23db869f You have the equation set up correctly in an earlier thread. Do you not know how you did it?

OpenStudy (anonymous):

NO

OpenStudy (anonymous):

30653.86 IS THE WRONG ANSWER I HAD THAT AND MY PROFESSER SAID IT WAS INCORRECT ANYONE ELSE HAS AN ANSWER THANKS

OpenStudy (anonymous):

Thanks for yelling. I see what might be the problem, but not I'm not too worried about it any more.

OpenStudy (anonymous):

dMANCINE I was not intentionally yelling just always keep my caps on I hope you don' t wy about it very tacky comment

OpenStudy (anonymous):

Now I'm tacky.

OpenStudy (anonymous):

Yes it was an error on my part for using caps but to make a comment like that in your tone was just as tacky as my caps.